OpenStudy (anonymous):

Write the following as simplified fraction 1) 12.45 2) 18 2/5% 3) 25.55%

OpenStudy (anonymous):

1.\[12 \frac{ 9 }{ 20 }\] 2.\[\frac{ 23 }{ 125}\] 3.\[\frac{ 511 }{ 2000 }\]

OpenStudy (anonymous):

1; (9/20) = 0.45